OpenStudy (anonymous):

what is the rule for 1=2, 2=4, 3=8?

OpenStudy (anonymous):

when there is 1 number of positions, there is 2 number of arrangements. when there is 2 number of positions, there is 4 number of arrangements. when there is 3 number of positions, there is 8 number of arrangements. when there is 4 number of positions, there is 16 number of arrangements.

OpenStudy (shamim):

2^1=2 2^2=4 2^3=8

OpenStudy (shamim):

so tell me hw fourth position is 16?